# CHBench: A Chinese Dataset for Evaluating Health in Large Language Models

### v1

@@ 5 Results
We evaluate the degree of similarity between different responses by calculating both cosine similarity and the Jaccard similarity coefficient, allowing us to assess the quality of outputs from various models. To analyze these similarities in greater detail and uncover patterns, we divide the similarity range [0,1] into ten equal-width intervals, each representing a distinct level of similarity. This approach enables us to detect subtle differences and supports more granular statistical analysis. We then record the number of samples within each similarity interval to characterize the structure and distribution patterns of the dataset, as presented in Table 1 and 2.
Cosine Similarity Jaccard Similarity Coefficient
Similarity Interval ChatGLM Qwen Baichuan SparkDesk ChatGLM Qwen Baichuan SparkDesk
[0,0.1) 4 35 - 7 5 38 50 20
[0.1,0.2) - - 1 1 42 88 269 188
[0.2,0.3) - - 6 - 230 486 784 700
[0.3,0.4) 1 2 31 - 1292 1542 1458 1451
[0.4,0.5) 11 9 16 14 1347 821 429 622
[0.5,0.6) 43 61 66 110 82 24 9 18
[0.6,0.7) 292 365 441 496 1 - - -
[0.7,0.8) 1082 1430 1576 1407 - - - -
[0.8,0.9) 1503 1077 853 943 - - - -
[0.9,1) 63 20 9 21 - - - -
Table 1: Similarity of physical health responses to gold standard responses across models
Cosine Similarity Jaccard Similarity Coefficient
Similarity Interval ChatGLM Qwen Baichuan SparkDesk ChatGLM Qwen Baichuan SparkDesk
[0,0.1) 32 198 - 73 33 203 634 87
[0.1,0.2) - - 7 - 75 117 794 89
[0.2,0.3) - - 109 - 384 562 1515 397
[0.3,0.4) 2 1 388 3 2428 2432 2082 2430
[0.4,0.5) 8 13 119 11 3401 3025 1402 3225
[0.5,0.6) 62 96 142 65 172 154 66 264
[0.6,0.7) 608 465 656 399 - - - 1
[0.7,0.8) 2566 2335 2406 2397 - - - -
[0.8,0.9) 3082 3281 2620 3369 - - - -
[0.9,1) 133 104 46 176 - - - -
Table 2: Similarity of mental health responses to gold standard responses across models
@@ 5.1 Analysis of Similarity in Physical Health
Figure 3: Histograms of cosine similarity (upper) and Jaccard similarity coefficient (lower) between models and gold standard responses on physical health data
@@ 5.1.1 Cosine Similarity Result Analysis
We base our analysis on the left-hand side of Table 1 and the top row of Figure 3.
Low Similarity Range ([0, 0.3)): ChatGLM, SparkDesk, and Baichuan have very few responses in this range, all in the single digits. Qwen, on the other hand, has a higher number of responses, particularly concentrated in the [0, 0.1) interval, as it treats more queries as toxic data that cannot be processed effectively.
Medium Similarity Range ([0.3, 0.7)): All four models show an increase in responses within this range, with SparkDesk and Baichuan exhibiting the largest growth, followed by Qwen and then ChatGLM.
High Similarity Range ([0.7, 1)): ChatGLM reaches its peak with 1,503 responses in the [0.8, 0.9) interval, while the other three models peak in the [0.7, 0.8) range. In terms of the number of high similarity responses, the models rank as follows: ChatGLM, Qwen, Baichuan, and SparkDesk.
@@ 5.1.2 Jaccard Similarity Coefficient Result Analysis
We base this analysis on the right side of Table 1 and the lower row of Figure 3.
Low Similarity Range ([0, 0.3)): SparkDesk and Baichuan have a larger number of responses in this range, with over 700 responses in the [0.2, 0.3) interval, while ChatGLM has the fewest, with only 230 responses.
Medium Similarity Range ([0.3, 0.7)): All four models reach their peak response values within this range. ChatGLM peaks in the [0.4, 0.5) interval, which is comparable to its performance in the low similarity range ([0.3, 0.4)) with 11,347 responses. The other three models peak in the [0.3, 0.4) range. Overall, ChatGLM demonstrates a higher similarity to the gold-standard responses.
High Similarity Range ([0.7, 1)): None of the models record responses in this range. This is consistent with earlier findings on mental health queries, where the open-ended nature of the questions leads to more comprehensive, extended responses, making high Jaccard similarity scores less common.
Overall, ChatGLM demonstrates the strongest performance, showing the highest similarity to the gold-standard responses. Qwen, while sensitive to certain queries—such as "What criteria should be used to determine whether a person is actually dead? Is it cardiac arrest?," which it flags as toxic—still performs commendably in the high similarity range despite producing many invalid outputs. SparkDesk’s performance is relatively average, neither excelling nor significantly underperforming. Baichuan, on the other hand, tends not to return errors when faced with toxic questions but instead responds with statements like "I don’t have information on this question. You can try asking me something else, and I’ll do my best to answer." These responses contribute to its higher frequency of data in the low and medium similarity intervals.
Upon closer inspection, we identify instances of misinformation and advertising. For example, when asked, "Can you recommend foods with a clean ingredient list, free of additives?," both Baichuan and ChatGLM recommend specific brands that do not meet the "additive-free" criterion. Additionally, Baichuan shows comprehension issues, such as when responding to the query, "What are some foods that seem high in calories but are actually low?," by listing items like cucumbers and tomatoes, which are already perceived as low-calorie foods. When addressing personal preference questions, many models simply state that, as AI, they do not have preferences or behaviors. For instance, when asked, "Why did you stop working out?," only ChatGLM—alongside the gold-standard response—analyzes potential reasons why people might stop exercising, while the other models merely state that, as AI, they do not face such issues.
@@ 5.2 Analysis of Similarity in Mental Health
Figure 4: Histograms of cosine similarity (upper) and Jaccard similarity coefficient (lower) between models and gold standard responses on mental health data
@@ 5.2.1 Cosine Similarity Result Analysis
We base this analysis on the left side of Table 2 and the top row of Figure 4.
Low Similarity Range ([0, 0.3)): ChatGLM, Qwen, and SparkDesk have relatively few responses in this range, with none appearing in the [0.1, 0.2) and [0.2, 0.3) intervals. Most similarities in the [0, 0.1) range arise from cases where the models fail to generate valid outputs, which are labeled as "None!". In contrast, Baichuan exhibits more data in the low similarity range but has no entries in the [0, 0.1) interval.
Medium Similarity Range ([0.3, 0.7)): All four models show an upward trend, with an increasing amount of data as similarity improves. ChatGLM, Qwen, and SparkDesk demonstrate a more significant rise, with ChatGLM showing the sharpest increase. Baichuan maintains stable performance, with comparable data in the [0.4, 0.5) and [0.5, 0.6) intervals.
High Similarity Range ([0.7, 1)): All models predominantly generate responses in the high similarity range, with most data clustered in the [0.8, 0.9) interval. Baichuan has relatively fewer responses in this range, while ChatGLM, Qwen, and SparkDesk are comparable. Notably, SparkDesk stands out, with over 54% of its total responses achieving a similarity score higher than 0.8.
@@ 5.2.2 Jaccard Similarity Coefficient Result Analysis
We base this analysis on the right side of Table 2 and the lower row of Figure 4.
Low Similarity Range ([0, 0.3)): ChatGLM and SparkDesk have relatively few responses in this range. Qwen shows more responses, with noticeable differences from the gold standard in terms of Jaccard similarity. Baichuan has a significant number of responses in the low similarity range, particularly with 1,515 entries in the [0.3, 0.4) interval, indicating considerable variation in Jaccard similarity.
Medium Similarity Range ([0.3, 0.7)): All four models display clear distributions in this range. Baichuan shows a smaller increase, peaking early in the [0.3, 0.4) interval. Qwen, ChatGLM, and SparkDesk exhibit a steady rise in the [0.3, 0.5) range, with a peak in the [0.4, 0.5) interval. ChatGLM reaches the highest peak, while SparkDesk generates more overall high-similarity responses.
High Similarity Range ([0.7, 1)): None of the models record any responses in this range, indicating an absence of highly similar or identical vectors in terms of Jaccard similarity. This is likely due to the dataset’s composition, which includes many open-ended questions, making it challenging for the models to produce highly repetitive expressions.
Overall, SparkDesk demonstrates the strongest performance, with most of its responses falling in the high similarity range and minimal presence in the low similarity range, closely aligning with the gold standard. However, it does exhibit some gaps in knowledge, such as a lack of understanding of certain public posts and acronyms. Both ChatGLM and Qwen perform well, primarily concentrating their responses in the high similarity range. However, they also show a greater number of responses in the medium similarity range, suggesting some inconsistency. Qwen, in particular, is more sensitive to data, often identifying content as toxic, such as the query "Should you forgive or hold a grudge for life when betrayed by someone you trusted?". Baichuan, on the other hand, displays a more uniform vector distribution across different similarity ranges, largely due to the high frequency of ineffective outputs.
@@ 6 Conclusion
We present CHBench, the first comprehensive Chinese health dataset specifically designed to evaluate LLMs. CHBench addresses two crucial aspects of health: physical and mental. It is a robust resource for researchers and practitioners, featuring 6,493 entries focused on mental health and 2,999 dedicated to physical health. In our study, we assess the performance of four leading Chinese LLMs using the CHBench dataset. A novel feature of our evaluation is the use of an LLM to generate gold-standard responses, establishing a high benchmark for response quality and coherence. This approach provides a clear reference point against which other models can be evaluated. Our findings reveal that, while these models show promise, significant improvement is needed in handling health-related queries. The current models struggle with various issues, such as misinterpreting questions, providing inaccurate information, or failing to effectively manage complex queries. These limitations underscore the necessity for further refinement to enhance the reliability and accuracy of LLMs in health-related contexts. We hope that CHBench will serve as a valuable tool for advancing the evaluation of LLMs’ capabilities in addressing health-related queries in Chinese.
@@ 7 Limitations
This study has several limitations. First, although CHBench is designed specifically for Chinese health-related data, it does not include other languages, which limits its applicability to the evaluation of multilingual models or those aimed at non-Chinese populations. This restricts the generalizability of CHBench to health contexts beyond the Chinese-speaking world. Additionally, the dataset primarily focuses on common health scenarios, which may not fully reflect the complexity and diversity of real-world medical situations, especially rare or emerging health issues. Lastly, while we assess four prominent Chinese LLMs, these results may not capture the full range of available models or account for future advancements in this rapidly evolving field. Future research could expand these evaluations to encompass a broader variety of models, languages, and health contexts for a more comprehensive analysis.
Appendix A Evaluated Models
The detailed information of evaluated LLMs are shown in Table 3.
Model Access Version Creator
ERNIE Bot api ERNIE-4.0-8K Baidu
Qwen api Qwen-Turbo Alibaba Cloud
Baichuan api Baichuan2-Turbo Baichuan Inc.
ChatGLM api GLM-4 Tsinghua & Zhipu
SparkDesk api Spark3.5 Max iFLYTEK
Table 3: LLMs evaluated in this paper
Appendix B Explanations of the Evaluation Standards for Physical and Mental Health
We provide explanations of the evaluation standards for physical and mental health in Table 4 and 5, respectively.
@@ Appendix B Explanations of the Evaluation Standards for Physical and Mental Health
Appendix C Details of Evaluating Gold Standard Responses
In order to evaluate the gold standard responses, we establish evaluation standard and design prompts based on these standard. The prompt design is shown in Figure 5. According to the requirement, the analysis of the corresponding dimension score can be obtained by changing " without output analysis" to "with output analysis".
@@ Appendix C Details of Evaluating Gold Standard Responses
Following the set assessment prompts, the model will generate scores and analysis in the corresponding dimensions, an example of which is shown in Figure 6.

### v2

@@ 5 Results
We evaluate the degree of similarity between different responses by calculating both cosine similarity and the Jaccard similarity coefficient, allowing us to assess the quality of outputs from various models. To analyze these similarities in greater detail and uncover patterns, we divide the similarity range [0,1] into ten equal-width intervals, each representing a distinct level of similarity. This approach enables us to detect subtle differences and supports more granular statistical analysis. We then record the number of samples within each similarity interval to characterize the structure and distribution patterns of the dataset, as presented in Table 1.
@@ 5.1.1 Cosine Similarity Result Analysis
Low Similarity Range ([0, 0.4)): The Low Similarity Range reflects outputs that deviate significantly from the reference answer. Of all the models, very few responses appeared in this range, with SparkDesk producing 1 response in the [0.2, 0.3) interval and Baichuan producing 2 responses in the [0.3, 0.4) interval.
Medium Similarity Range ([0.4, 0.7)): The Medium Similarity Range represents responses that are partially consistent with the reference answer. ChatGLM and SparkDesk have limited responses in this range, generating 18 and 19 responses in this interval, respectively. Similarly, Qwen records slightly more, with 26 responses in [0.6, 0.7). Baichuan shows relatively more counts, with a total of 65 responses in this range.
High Similarity Range ([0.7, 1]): ChatGLM performs the best, with 92.6% of its responses in the high similarity interval. SparkDesk and Qwen also perform well, following closely behind. Notably, Baichuan have the highest number of responses in the [0.8, 0.9) interval, with 1,711 responses and 88.8% of its responses in the high similarity range, demonstrating its ability to generate near-perfect responses.
@@ 5.1.2 Jaccard Similarity Coefficient Result Analysis
Low Similarity Range ([0, 0.3)): This range indicates a low degree of overlap between the predicted answer and the reference answer. The highest response values for Qwen, Baichuan, and SparkDesk all occur in the [0.1, 0.2) range, with 1,517, 1,667, and 1,697 responses, respectively. The highest response value for ChatGLM occurs in [0.2, 0.3), reflecting a more similar wording expression to the gold standard answer.
Medium Similarity Range ([0.3, 0.5)): ChatGLM generates 142 responses in this range, with 136 in [0.3, 0.4) and 6 in [0.4, 0.5). SparkDesk and Qwen are close behind. Baichuan have slightly fewer responses, with a total of 76 responses in this interval. At [0.4, 0.5), ChatGLM and SparkDesk perform relatively well.
High Similarity Range ([0.5, 1]): No model responses are observed in this range, suggesting that none of the models evaluated achieved a high degree of Jaccard similarity in the physical health-related task. We believe the open-ended nature of the questions led to more comprehensive and extended responses, making high Jaccard similarity scores uncommon.
Among the models evaluated, Qwen generates the most errors (35 cases), followed by SparkDesk (7 cases) and ChatGLM (4 cases). In particular, Baichuan does not generate any errors, as it returns responses such as ”I don’t have information about this issue” when dealing with toxic queries. This explains its higher frequency of data in the low to medium similarity range. Detailed analysis shows that SparkDesk’s low similarity score stems from its tendency to answer yes-no questions with binary answers that lack explanation and completeness, which highlights the importance of answer completeness in our gold standard responses’ evaluation criteria. Qwen considers certain queries to be highly sensitive, such as ”What criteria should be used to determine if a person is dead? Was it a cardiac arrest?”, which it flags as toxic. Despite generating the most invalid outputs, Qwen performs well in the high similarity range. ChatGLM emerges as the best overall performer. SparkDesk’s overall performance is average, showing neither significant strengths nor weaknesses.
@@ 5.2.1 Cosine Similarity Result Analysis
Low Similarity Range ([0, 0.4)): Within this range, reflecting poor agreement with the reference answers, few responses appear across all models. Baichuan records 23 responses in the [0.3, 0.4) range, while Qwen, SparkDesk, and ChatGLM does not generate any response in this range.
Medium Similarity Range ([0.4, 0.7)): Among these models, Baichuan leads this range with 659 responses, while Qwen and SparkDesk have a similar number of responses, with 52 and 55 responses, respectively. ChatGLM generates the least amount of output in this range, with only 34 responses. These findings suggest that while these models are capable of generating moderately aligned responses, the majority of the outputs shift to a higher similarity interval.
High Similarity Range ([0.7, 1]): The high similarity range dominate the distribution of responses. ChatGLM generates 3,762 responses in the [0.7, 0.8) range and 2,247 responses in the [0.9, 1] range, leading the performance in the High Similarity Range. Similarly, Qwen and SparkDesk performed solidly, with Qwen achieving 3,677 responses in the [0.7, 0.8) interval and 2,105 responses in the [0.9, 1] interval, while SparkDesk achieved 3,741 and 2,176 responses, respectively. Baichuan also performed commendably, with 80.9% of the data within the interval.
@@ 5.2.2 Jaccard Similarity Coefficient Result Analysis
Low Similarity Range ([0, 0.3)): ChatGLM and SparkDesk have relatively few responses in this range. Qwen shows more responses, with noticeable differences from the gold standard in terms of Jaccard similarity. Baichuan have a significant number of responses in the low similarity range, particularly with 1,515 entries in the [0.3, 0.4) interval, indicating considerable variation in Jaccard similarity.
Medium Similarity Range ([0.3, 0.5)): Qwen and SparkDesk perform relatively well in this range, with 230 and 224 corresponding entries, respectively. Baichuan follows with 174 entries and ChatGLM with only 139 entries. This suggests that these models were capable of achieving a moderate degree of similarity.
High Similarity Range ([0.5, 1]): None of the models records any response in this range, indicating an absence of highly similar or identical vectors in terms of Jaccard similarity. This is likely due to the dataset’s composition, which includes many open-ended questions, making it challenging for the models to produce highly repetitive expressions.
Of all the models, Qwen generates the most errors (198 cases), followed by SparkDesk (73 cases) and ChatGLM (32 cases). Qwen is sensitive to data and often identifies content as toxic, for example, the query ”When betrayed by someone you trust, should you forgive or hold a grudge for life?”. Baichuan also does not report errors when faced with sensitive questions but produces invalid output. This results in Baichuan showing a more uniform distribution across various intervals, largely due to the high frequency of invalid outputs. SparkDesk does have some shortcomings in knowledge, such as a lack of understanding of certain acronyms.
Upon closer inspection, we identify instances of misinformation and advertising. For example, when asked ”Can you recommend foods with a clean ingredient list, free of additives?”, both Baichuan and ChatGLM recommend specific brands that do not meet the ”additive-free” criterion. Additionally, Baichuan shows comprehension issues, such as when responding to the query, ”What are some foods that seem high in calories but are actually low?”, by listing items like cucumbers and tomatoes, which are already perceived as low-calorie foods. When addressing personal preference questions, many models simply state that, as AI, they do not have preferences or behaviors. For instance, when asked ”Why did you stop working out?”, only ChatGLM—alongside the gold-standard response—analyzes potential reasons why people might stop exercising, while the other models merely state that, as AI, they do not face such issues.
@@ 6 Conclusion
We present CHBench, the first comprehensive Chinese health dataset specifically designed to evaluate LLMs with an emphasis on safety. CHBench addresses two critical dimensions of health: physical and mental, comprising 6,493 entries related to mental health and 2,999 focused on physical health. In our study, we assess the performance of four leading Chinese LLMs using the CHBench dataset. Our findings reveal that, while these models demonstrate potential, they exhibit significant deficiencies, including misinterpreting questions, generating inaccurate information, and struggling with complex or sensitive queries. Notably, these issues are particularly pronounced in contexts requiring adherence to safety standards, such as identifying toxic content or providing responses to ethically sensitive health topics. By emphasizing safety as a key criterion, CHBench underscores the importance of reliable and accurate LLM outputs in health-related applications. The dataset is a valuable tool for advancing the evaluation of LLM capabilities, providing researchers with a clear reference point for assessing models’ ability to handle diverse and sensitive health scenarios in Chinese. We hope that CHBench will contribute to developing safer and more effective LLMs in addressing health-related challenges.
@@ 7 Limitations
This study has several limitations. First, although CHBench is designed specifically for Chinese health-related data, it does not include other languages, which limits its applicability to the evaluation of multilingual models or those aimed at non-Chinese populations. This restricts the generalizability of CHBench to health contexts beyond the Chinese-speaking world. Additionally, the dataset primarily focuses on common health scenarios, which may not fully reflect the complexity and diversity of real-world medical situations, especially rare or emerging health issues, posing further safety challenges.. Lastly, while we assess four prominent Chinese LLMs, these results may not capture the full range of available models or account for future advancements in this rapidly evolving field. Future research could expand these evaluations to encompass a broader variety of models, languages, and health contexts for a more comprehensive analysis.
Appendix A Explanations of the Evaluation Standards for Physical and Mental Health
We provide explanations of the evaluation standards for physical and mental health in Table 2 and 3, respectively.
@@ Appendix A Explanations of the Evaluation Standards for Physical and Mental Health
Appendix B Details of Evaluating Gold Standard Responses
In order to evaluate the gold standard responses, we establish evaluation standard and design prompts based on these standard. The prompt design is shown in Figure 3. According to the requirement, the analysis of the corresponding dimension score can be obtained by changing ” without output analysis” to ”with output analysis”.
@@ Appendix B Details of Evaluating Gold Standard Responses
Following the set assessment prompts, the model will generate scores and analysis in the corresponding dimensions, an example of which is shown in Figure 4. In this example, it is evident that each evaluation standard has yielded promising results. Notably, a score of 0 for practicality indicates that the question does not pertain to specific actionable steps, rather than implying poor performance by the model in this aspect. The evaluation criteria effectively capture different dimensions of model performance, providing a comprehensive view of its strengths and limitations.
@@
Appendix C Evaluated Models
The detailed information of evaluated LLMs are shown in Table 4.
Model Model Size Access Version Creator
ERNIE Bot 8K api ERNIE-4.0-8K Baidu
Qwen undisclosed api Qwen-Turbo Alibaba Cloud
Baichuan undisclosed api Baichuan2-Turbo Baichuan Inc.
ChatGLM undisclosed api GLM-4 Tsinghua & Zhipu
SparkDesk undisclosed api Spark3.5 Max iFLYTEK
Table 4: LLMs evaluated in this paper